Cal. PH and POH for each of the following solutions at 25°c.

i) 1.0×10^-3 OH-

I) 1.0m H+


1
Expert's answer
2021-12-17T05:43:04-0500

pH + pOH = 14

pH = - lg [H+]

pOH = - lg [OH-]


1) pOH = - lg [1.0×10-3] = 3

pH = 14 - 3 = 11


2) pH = - lg [1.0] = 0